Write the equation in standard form using only integers
(1/2)x-9=0

Answer by checkley75(3666)

(1/2)X-9=0
X/2=9 NOW CROSS MULTIPLY:
X=2*9
X=18
PROOF
(1/2)18-9=0
9-9=0
0=0